OpenStudy (anonymous):

Determine if the following function is odd, even, or neither. f(x) = -5x2 + 3x - 1

OpenStudy (anonymous):

what is the value of x?

OpenStudy (anonymous):

This function is neither. A function is odd if f(-x) = -f(x); the function will be even if f(-x) = f(x).

OpenStudy (anonymous):

the value of x is not important here. all you need to look at is the polynomial this function is even because it has y axis symetry

OpenStudy (anonymous):

it would look something like this 

OpenStudy (anonymous):

f(-x) = 5x2+3x-1 f(-x) =5(-x)+3(-x)-1 f(-x)= -5x2-3x-1 f(-x)= -(5x2-3x-1) f(-x) = -f(x) so it is odd function